I need help solving this rational equation t+5/t-2=t-2/t+4
Answer by nadim(2)

ok,follow me,
t+5/t-2=t-2/t+4
or,5/t+2/t=4+2
or,(5+2)/t=6
or,7/t=6
or,t=7/6